# sample.bib

### v1

@@ 2 Algorithm
Let $\mathcal{D}=\{z_{1},...,z_{n}\}$ be a training dataset of $n$ data points drawn independently and identically distributed from the sample space $\mathcal{Z}$, and let $\Theta$ be the model parameter space. Let $A:\mathcal{Z}^{n}\to\Theta$ be a (randomized) learning algorithm that trains on $\mathcal{D}$ and outputs a model with weight parameters $\theta\in\Theta$, where $\Theta$ is the (potentially infinite) space of possible model weights. Typically, the goal of a learning algorithm is to minimize $f_{\mathcal{D}}(\theta)$, defined as the empirical loss on $\mathcal{D}$ as follows
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta)$
@@
Let us “unlearn” or remove the influence of a subset of data $Z\subset\mathcal{D}$ from the output of the learning algorithm $A(\mathcal{D})$. Let $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$, and we denote by $U(A(\mathcal{D}),\mathcal{D},Z)$ the output of an unlearning algorithm $U$. The goal of the unlearning algorithm is to output a model parameter that is probabilistically indistinguishable from the output of $A(\mathcal{D^{\prime}})$. This is formalized in the concept of $(\epsilon,\delta)$-indistinguishability, which is used in the differential privacy literature to characterize the influence of a data point on the model output [dwork2014algorithmic].
@@
[dwork2014algorithmic] [neel21a] Let $X$ and $Y$ be random variables over some domain $\Omega$. We say $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if for all $S\subseteq\Omega$ we have
@@
dataset $\mathcal{D}$, initial point $\theta_{0}\in\Theta$
@@
     $\theta_{t}=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})$triangleright Gradient descent steps
@@
$\tilde{\theta}=\theta_{T}+\xi$triangleright Gaussian perturbation
@@
Upon receiving an unlearning request, execute Algorithm 2
@@
dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T-K}$
@@
     $\theta^{\prime\prime}_{t}=\theta^{\prime\prime}_{t-1}-\eta\nabla f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{t-1})$triangleright Gradient descent steps
@@
$\tilde{\theta}^{\prime\prime}=\theta^{\prime\prime}_{K}+\xi$triangleright Gaussian perturbation
@@
When a model is trained without a checkpoint saved, we can still obtain a black-box unlearning algorithm by carefully “rewinding” the gradient descent training steps via the proximal point method, outlined in Algorithm 3. In standard gradient descent, we obtain $\theta_{t+1}$ from $\theta_{t}$ as follows,
$\theta_{t+1}=\theta_{t}-\eta\nabla f_{\mathcal{D}}(\theta_{t}).$
When we solve for previous gradient descent iterates, we obtain the following implicit equation:
@@
We note that the backward Euler update in (1) is distinct from a standard gradient ascent step, which is a forward Euler method. Instead, we can compute $\theta_{t}$ from $\theta_{t+1}$ by taking advantage of the connection between backward Euler for gradient flow and the proximal point method, an iterative algorithm for minimizing a convex function [Martinet_1970prox]. Let $g(\theta)$ denote a convex function. The proximal point method minimizes $g(\theta)$ by taking the proximal operator with parameter $\gamma$ of the previous iterate, defined as follows
@@
Although for our problem, $f$ is nonconvex, sufficient regularization can produce a convex and globally tractable proximal point subproblem.
Lemma 2.3.
[drusvyatskiy2017proximalpointmethodrevisited] If $f(\theta)$ is continuously differentiable with $L$-Lipschitz gradient, $-f(\theta)+\frac{L}{2}||\theta||^{2}$ is convex.
Finally, by computing the proximal operator with respect to $-f(\theta)$, we can solve the implicit gradient ascent step.
@@
Proof.
Suppose $\theta^{*}=prox_{-f,\eta}(\theta_{t+1})=\arg\min_{x}\{-f(x)+\frac{1}{2\eta}||x-\theta_{t+1}||^{2}\}$, which is well-defined due to Lemma 2.3. Then the gradient of the objective function is zero at $\theta^{*}$ and thus
$-\nabla f(\theta^{*})+\frac{1}{\eta}(\theta^{*}-\theta_{t+1})=0,$
$\theta^{*}=\theta_{t+1}+\eta\nabla f(\theta^{*}).$
@@
datasets $\mathcal{D}$, model checkpoint $\theta_{T}$
@@
     $\theta^{\prime\prime}_{t}=\arg\min_{x}\{-f_{\mathcal{D}}(x)+\frac{1}{2\eta}||x-\theta^{\prime\prime}_{t-1}||^{2}\}$triangleright Proximal point subproblem
@@
If $\eta<\frac{1}{L}$, then due to convexity, we can solve the proximal point subproblem easily via gradient descent or Newton’s method. Computationally, when $K\ll T$, the algorithm is comparable to other second-order unlearning methods that require a single Newton step. In addition, we only need to compute the model checkpoint once prior to unlearning requests, so this computation can be considered “offline” [izzo21a] [qiao2024efficientgeneralizablecertifiedunlearning].
Our approach is inspired by first-order methods such as [neel21a] [chien2024langevin] [chien2024stochastic]. In the aforementioned works, the learning algorithm $A$ is some variant of gradient descent on $f_{\mathcal{D}}$, and the unlearning algorithm $U$ is first initialized at the output weights of $A$ before proceeding with some number of additional gradient descent iterations on $f_{\mathcal{D^{\prime}}}$. Intuitively, because these works assume a unique minimizer or limiting distribution that attracts all gradient descent trajectories, the unlearning iterates will converge to the output $A(\mathcal{D^{\prime}})$. However, because gradient descent on nonconvex functions may converge to flat regions or saddle points, there is no guarantee that additional iterates after the learning process will make progress towards the destination of the retraining trajectory. Therefore, checkpointing $K$ iterates before the final timestep backtracks the algorithm to a steeper region of the loss landscape and draws the trajectory closer to the retraining trajectory.
@@ 3 Analyses
In the following theorem, we establish the unlearning and performance guarantees for our algorithm on nonconvex functions. For nonconvex functions, gradient descent might converge to local minima or saddle points, so we measure the performance by the average of the gradient norm over iterates. In differential privacy, gradient norm is a common performance metric for algorithms on nonconvex functions.
@@
Let $\epsilon,\delta>0$ be fixed. Suppose for all $z\in\mathcal{Z}$, the loss function $f_{z}$ is $L$-Lipschitz smooth and the gradient is uniformly bounded by some constant $G$ so that $\lVert\nabla f_{z}(\theta)\rVert<G$ for all $\theta\in\Theta$. Let $D$ denote the original dataset of size $n$, let $Z\subset\mathcal{D}$ denote the unlearned dataset of size $m$, and let $\mathcal{D}^{\prime}=\mathcal{D}\backslash Z$ denote the retain set. Let the learning algorithm $A$ be initialized at $\theta_{0}$ and run for $T$ iterations with step-size $\eta\leq\min\{\frac{1}{L},\frac{n}{2(n-m)L}\}$. Let the standard deviation $\sigma$ of the Gaussian noise be defined as
@@
$\frac{1}{T}\Big{[}\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\sum_{t=0}^{K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}+\mathbb{E}[||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}]\Big{]}\leq$ $\frac{2n(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))}{T\eta(n-m)}$ (3)
$+\frac{T-K-1}{T}(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})$
$\leq$ $O(\frac{n}{T(n-m)})+O(\frac{(T-K-1)m}{T(n-m)})$
@@
Equation (3) states that the average of the gradient norm squared of the initial $T-K$ learning iterates, the $K-1$ unlearning iterates after, and the last perturbed iterate $\tilde{\theta}^{\prime\prime}$ decreases with increasing $T$ and $n$, showing that the algorithm converges to a stationary point with small gradient norm.
The proof of Theorem 3.1 is provided in Appendix B. The proof relies on carefully tracking the distance between the unlearning iterates and gradient descent iterates on $f_{\mathcal{D^{\prime}}}$. Like prior work in certified unlearning, our analysis relies on the Gaussian mechanism for differential privacy [dwork2014algorithmic], which implies that as long as the distance between the trajectories is bounded, we can add a sufficient amount of Gaussian noise to make the algorithm outputs ($\epsilon$, $\delta$)-indistinguishable. This allows us to determine the noise level $\sigma$ required to achieve unlearning. For the utility guarantees, we leverage the fact that gradient descent steps on $f_{\mathcal{D}}$ also make progress on $f_{\mathcal{D^{\prime}}}$ to obtain bounds that only depend on problem parameters and the initialization $\theta_{0}$ and do not depend on points along the training trajectory.
Theorem 3.1 underscores the “privacy-utility-complexity” tradeoff between our measure of unlearning, $\epsilon$, noise, $\sigma$, and the number of unlearning iterations, $K$. By construction, $K<T$, so our algorithm is more efficient than retraining. We can pick larger $K$ such that the noise required is arbitrarily small at the expense of computation, and when $K=T$ our algorithm amounts to a noiseless full retrain. Moreover, the standard deviation $\sigma$ inversely scales with the size of the dataset $n$, implying that unlearning on larger datasets require less noise. In contrast, [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] do not feature such data-dependent guarantees.
Remark.
The guarantee in Theorem 3.1 applies to unlearning a batch of $m$ data samples. Our algorithm also accommodates sequential unlearning requests. If, after unlearning $m$ points, an additional $k$ unlearning requests arrive, we can simply call the unlearning algorithm $M$ on the new retain set of size $n-m-k$. Notably, if the total number of unlearned data increases while $\sigma$ and $K$ stay constant, our unlearning guarantee worsens, which is consistent with other unlearning algorithms [zhang2024towards][chien2024langevin][Guo].
@@
[karimiPL] For some function $f$, suppose it attains a global minimum value $f^{*}$. Then $f$ satisfies the PL inequality if for some $\mu>0$ and all $x$,
@@
Since PL functions have a continuous basin of global minima, we can obtain empirical risk bounds. Corollary 3.3 states that the empirical risk converges linearly with both $T$ and $K$. The proof is provided in Appendix B.1.
@@
in order to determine how well the model will generalize on unseen test data. We leverage results from [lei2021sharper] that relate the on-average stability bounds of algorithms on PL functions to their excess population risk.
@@
$\mathbb{E}[F(\tilde{\theta}^{\prime\prime})]-F^{*}\leq$ $L\sqrt{d}\sigma+\frac{2G^{2}}{(n-m)\mu}$
$+\frac{L}{2\mu}\left((1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}\right)$

### v2

@@ 2 Algorithm
Let $\mathcal{D}=\{z_{1},...,z_{n}\}$ be a training dataset of $n$ data points drawn independently and identically distributed from the sample space $\mathcal{Z}$, and let $\Theta$ be the model parameter space. Let $A:\mathcal{Z}^{n}\to\Theta$ be a (randomized) learning algorithm that trains on $\mathcal{D}$ and outputs a model with weight parameters $\theta\in\Theta$, where $\Theta$ is the (potentially infinite) space of model weights. Typically, the goal of a learning algorithm is to minimize $f_{\mathcal{D}}(\theta)$, the empirical loss on $\mathcal{D}$, defined as follows
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta),$
@@
Let us “unlearn” or remove the influence of a subset of data $Z\subset\mathcal{D}$ from the output of the learning algorithm $A(\mathcal{D})$. Let $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$, and we denote by $U(A(\mathcal{D}),\mathcal{D},Z)$ the output of an unlearning algorithm $U$. The goal of the unlearning algorithm is to output a model parameter that is probabilistically indistinguishable from the output of $A(\mathcal{D^{\prime}})$. This is formalized in the concept of $(\epsilon,\delta)$-indistinguishability, which is used in the DP literature to characterize the influence of a data point on the model output (Dwork & Roth, 2014).
@@
(Dwork & Roth, 2014; Neel et al., 2021) Let $X$ and $Y$ be random variables over some domain $\Omega$. We say $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if for all $S\subseteq\Omega$,
@@
0:  dataset $\mathcal{D}$, initial point $\theta_{0}\in\Theta$
@@
     $\theta_{t}=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})$Gradient descent steps}
@@
  $\tilde{\theta}=\theta_{T}+\xi$                      {Gaussian perturbation}
@@
  Upon receiving an unlearning request, call Algorithm 2
@@
0:  dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T-K}$
@@
     $\theta^{\prime\prime}_{t}=\theta^{\prime\prime}_{t-1}-\eta\nabla f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{t-1})$Gradient descent steps}
@@
  $\tilde{\theta}^{\prime\prime}=\theta^{\prime\prime}_{K}+\xi$                      {Gaussian perturbation}
@@
When a model is trained without a checkpoint saved, we can still obtain a black-box unlearning algorithm by carefully “rewinding” the gradient descent training steps via the proximal point method, outlined in Algorithm 3. We can solve for previous gradient descent iterates through the following implicit equation (2):
$\theta_{t+1}$ $=\theta_{t}-\eta\nabla f_{\mathcal{D}}(\theta_{t})$ (1)
@@
The backward Euler update in (2) is distinct from a standard gradient ascent step, which is a forward Euler method. Instead, we compute $\theta_{t}$ from $\theta_{t+1}$ by taking advantage of the connection between backward Euler for gradient flow and the proximal point method, an iterative algorithm for minimizing a convex function (Martinet, 1970). Let $g(\theta)$ denote a convex function. The proximal point method minimizes $g(\theta)$ by taking the proximal operator with parameter $\gamma$ of the previous iterate, defined as follows
@@
Although for our problem, $f$ is nonconvex, adding sufficient regularization produces a convex and globally tractable proximal point subproblem, stated in Lemma A.1. Therefore, by computing the proximal operator with respect to $-f(\theta)$, we can solve the implicit gradient ascent step.
@@
The proof of Lemma 2.3 is provided in Appendix A.
@@
0:  datasets $\mathcal{D}$, model checkpoint $\theta_{T}$
@@
     $\theta^{\prime\prime}_{t}=\arg\min_{x}\{-f_{\mathcal{D}}(x)+\frac{1}{2\eta}||x-\theta^{\prime\prime}_{t-1}||^{2}\}$
@@
If $\eta<\frac{1}{L}$, then due to convexity, we can solve the proximal point subproblem easily via gradient descent or Newton’s method. Computationally, when $K\ll T$, the algorithm is comparable to other second-order unlearning methods that require a single Newton step. In addition, we only need to compute the model checkpoint once prior to unlearning requests, so this computation can be considered “offline” (Izzo et al., 2021; Qiao et al., 2024).
@@ 3 Analyses
In the following theorem, we establish the unlearning and performance guarantees for our algorithm on nonconvex functions. For nonconvex functions, gradient descent might converge to local minima or saddle points, so we measure the performance by the average of the gradient norm over iterates, a common DP performance metric for algorithms on nonconvex functions.
@@
Let $\epsilon,\delta$ be fixed such that $0<\epsilon\leq 1$ and $\delta>0$. Suppose for all $z\in\mathcal{Z}$, the loss function $f_{z}$ is $L$-Lipschitz smooth and the gradient is uniformly bounded by some constant $G$ so that $\lVert\nabla f_{z}(\theta)\rVert<G$ for all $\theta\in\Theta$. Let $\mathcal{D}$ denote the original dataset of size $n$, let $Z\subset\mathcal{D}$ denote the unlearned dataset of size $m$, and let $\mathcal{D}^{\prime}=\mathcal{D}\backslash Z$ denote the retained dataset. Let the learning algorithm $A$ be initialized at $\theta_{0}$ and run for $T$ iterations with step size $\eta\leq\min\{\frac{1}{L},\frac{n}{2(n-m)L}\}$. Let the standard deviation $\sigma$ of the Gaussian noise be defined as
@@
$\frac{\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\sum_{t=0}^{K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}}{T}\\ +\frac{\mathbb{E}[||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}]}{T}\leq O(\frac{n}{T(n-m)})+O(\frac{(T-K-1)m}{T(n-m)})$ (4)
@@
Corollary 3.2.
For fixed $\sigma$ and $\delta$, the dependence of $K$ on $\epsilon$ in (3), denoted as $K(\epsilon)$, is upper-bounded as follows:
$K(\epsilon)\leq\frac{\log\Big{(}(1+\frac{\eta Ln}{n-m})^{T}-\frac{\sigma Ln\epsilon}{2mG\sqrt{2\log(1.25/\delta)}}\Big{)}}{\log(1+\frac{\eta Ln}{n-m})}.$ (5)
Equation (4) states that the average of the gradient norm squared of the initial $T-K$ learning iterates, the $K-1$ unlearning iterates after, and the last perturbed iterate $\tilde{\theta}^{\prime\prime}$ decreases with increasing $T$ and $n$, indicating that the algorithm converges to a stationary point with small gradient norm. Corollary 3.2 provides an upper bound on $K(\epsilon)$, such that in practice we can choose $K$ equal to this bound to ensure the privacy guarantee is achieved. We show in Figure 1(c) that this bound is close to tight for $0<\epsilon\leq 1$ and real-world parameters.
The proof of Theorem 3.1 is provided in Appendix B. The proof of Corollary 3.2 is in Appendix B.3. The analysis relies on carefully tracking the distance between the unlearning iterates and gradient descent iterates on $f_{\mathcal{D^{\prime}}}$. Like prior work in certified unlearning, our analysis relies on the Gaussian mechanism for differential privacy (Dwork & Roth, 2014), which implies that as long as the distance between the trajectories is bounded, we can add a sufficient amount of Gaussian noise to make the algorithm outputs ($\epsilon$, $\delta$)-indistinguishable. We therefore can compute the noise level $\sigma$ required to achieve unlearning. For the utility guarantees, we leverage the fact that gradient descent steps on $f_{\mathcal{D}}$ also make progress on $f_{\mathcal{D^{\prime}}}$ to obtain bounds that only depend on problem parameters and the initialization $\theta_{0}$.
Theorem 3.1 underscores the “privacy-utility-complexity” tradeoff between our measure of unlearning, $\epsilon$, noise, $\sigma$, and the number of unlearning iterations, $K$. By construction, $K<T$, so our algorithm is more efficient than retraining. We can pick larger $K$ such that the noise required is arbitrarily small at the expense of computation, and when $K=T$ our algorithm amounts to a noiseless full retrain. Moreover, the standard deviation $\sigma$ inversely scales with the size of the dataset $n$, implying that unlearning on larger datasets require less noise. In contrast, (Zhang et al., 2024) and (Qiao et al., 2024) do not feature such data-dependent guarantees.
Remark 3.3.
Theorem 3.1 applies to unlearning a batch of $m$ data samples. Our algorithm also accommodates sequential unlearning requests. If, after unlearning $m$ points, an additional $k$ unlearning requests arrive, we simply call the unlearning algorithm $M$ on the new retained dataset of size $n-m-k$. Notably, if the total number of unlearned data increases while $\sigma$ and $K$ stay constant, our unlearning guarantee worsens, which is consistent with other results (Guo et al., 2019; Zhang et al., 2024; Chien et al., 2024a).
@@
(Karimi et al., 2016) For some function $f$, suppose it attains a global minimum value $f^{*}$. Then $f$ satisfies the PL inequality if for some $\mu>0$ and all $x$,
@@
Although PL functions may be nonconvex, they have a continuous basin of global minima, so we can obtain empirical risk bounds. Corollary 3.5 states that the empirical risk converges linearly with both $T$ and $K$. The proof is provided in Appendix B.1.
@@
$\mathbb{E}[f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})]-f_{\mathcal{D^{\prime}}}^{*}\leq L\sqrt{d}\sigma+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}\\ +(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})$
@@
in order to determine how well the model will generalize on unseen test data. We leverage results from (Lei & Ying, 2021) that relate the on-average stability bounds of algorithms on PL functions to their excess population risk.
@@
$\mathbb{E}[F(\tilde{\theta}^{\prime\prime})]-F^{*}\leq L\sqrt{d}\sigma+\frac{2G^{2}}{(n-m)\mu}\\ +\frac{L}{2\mu}(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})\\ +\frac{L}{2\mu}(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$